Capital University, a mid-size university located in a capital city has approximately 10,000 students studying in four departments. While the University’s goal is not to make a profit, it is also important that the University does not operate at a loss. The only source of revenue is tuition and government grants. The government gives the University $14,000 a year for each full-time student.

Faculty:

Arts

Science

Engineering

Business

Number of Courses

120

200

150

150

Number of Students

4,800

2,000

600

2,700

Other Information:

  • Each Course is only offered once each year
  • Each student is enrolled in 10 classes per year
  • Each faculty member is paid $120,000 per year and teaches 6 courses per year
  • Total student support costs are $139,992,700 per year. These cost include student services and administration. Each student generally requires an equal amount of support.
  • Building costs are $12,400,000 per year. Costs are equal for each class.
  • Other costs $17,000,000 (Science labs $10million, Business trading room $2million, Engineering labs $5million).

REQUIRED

1. Using Activity Based Costing, calculate how much tuition should be charged to students in each faculty.
Answer #1

A.ABC rate calculation

Costs Cost drive Cost pool Number of cost driver ABC rate
Faculty salary Number of faculty 103.333 120000(Directly given
Student support cost Number of students 139992700 10100 13860.67
Building costs Number of classes 12400000 101000 122.77

Note

1.Number of faculty

Total number of course offered = 120+200+150+150 =620

Each faculty teaching course per year = 6 courses

Number of faculty =620/6 =103.33

2.Total number of students =4800+2000+600+2700=10100

3.Total class per year =Each student is enrolled in 10 classes per year

=Total student*10=10100*10=101000

B) Cost allocation Statement

Cost Arts science Engineering Management
Faculty salary

(120000*(120/6)

2400000

(120000*200/6)

4000000

(120000*150/6)

3000000

(120000*150/6)

3000000

Student support cost

(13860.67*4800)

66531216

(13860.67*2000)

27721340

(13860.67*600)

8316402

(13860.67*2700)

37423809

Building costs

(122.77*48000)

5892960

(122.77*20000)

2455400

(122.77*6000)

736620

(122.77*27000)

3314790

Other costs 10000000 2000000 5000000
Total costs 74824176 44176740 14053022 48738599
Less:Grant income for each student

(4800*14000)

67200000

(2000*14000)

28000000

(600*14000)

8400000

(2700*14000)

37800000

Total fees to be charged 7624176 16176740 5653022 10938599
Fees to be charged per student(Total fees/Total number of students) 1588.37 8088.37 9421.7 4051.33